### v1

# Utility-Learning Tension in Self-Modifying Agents

Abstract
As systems trend toward superintelligence, a natural modeling premise is that agents can self-improve along every facet of their own design. We formalize this with a five-axis decomposition and a decision layer, separating incentives from learning behavior and analyzing axes in isolation. Our central result identifies and introduces a sharp utility–learning tension, the structural conflict in self-modifying systems whereby utility-driven changes that improve immediate or expected performance can also erode the statistical preconditions for reliable learning and generalization. Our findings show that distribution-free guarantees are preserved iff the policy-reachable model family is uniformly capacity-bounded; when capacity can grow without limit, utility-rational self-changes can render learnable tasks unlearnable. Under standard assumptions common in practice, these axes reduce to the same capacity criterion, yielding a single boundary for safe self-modification. Numerical experiments across several axes validate the theory by comparing destructive utility policies against our proposed two-gate policies that preserve learnability.
1 Introduction
Classical learning theory—from realizable and agnostic PAC to information-theoretic and computational analyses—rests on a tacit premise: the learning mechanism is architecturally invariant. Parameters may adapt, but the agent’s update rules, representational scaffolding, topology, computational substrate, and meta-reasoning are treated as fixed. As capabilities trend toward strong open-ended autonomy, however, it is increasingly realistic to assume that advanced agents will self-improve broadly, rewriting not just weights but the very mechanisms by which they learn.
Evidence for this shift already exists. Reinforcement learning and meta-learning instantiate constrained self-change (Sutton and Barto,, 2018; Finn et al.,, 2017; Rajeswaran et al.,, 2019; Hospedales et al.,, 2022), while open-ended pipelines iterate code edits and tools (Zhang et al.,, 2025). Decision-theoretic proposals investigate provably utility-improving modifications (Schmidhuber,, 2005), safety analyses document pathologies (Orseau and Ring,, 2011), and metagoal frameworks aim to stabilize goal evolution (Goertzel,, 2024). What remains underdeveloped is a learning-theoretic account of post-modification behavior: when do seemingly rational self-changes preserve the conditions under which learning is possible, and when do they destroy them? We prove a policy-level learnability boundary: distribution-free PAC guarantees are preserved iff the policy-reachable family has uniformly bounded capacity. A simple Two-Gate guardrail (validation margin $\tau$ + capacity cap $K(m)$) keeps trajectories on the safe side and yields a VC-rate oracle inequality.
Our Contributions.
  • Policy boundary (iff). Under standard i.i.d. assumptions, distribution-free PAC learnability is preserved under self-modification iff the policy-reachable family has uniformly bounded capacity (VC/pseudodim).
  • Axis reductions. Architectural and metacognitive edits reduce to induced hypothesis families; substrate changes matter only via the induced family. Hence the boundary depends solely on the supremum capacity of the reachable family.
  • Two-Gate guardrail. A computable accept/reject rule (validation improvement by margin $\tau$ + capacity cap $K(m)$) ensures monotone true-risk steps and an oracle inequality at VC rates for the final predictor.
2 Related Work
2.1 Decision-theoretic self-modification and safety
Gödel Machines give a proof-based framework for globally optimal self-modification under a utility function (Schmidhuber,, 2005). Safety analyses document pathologies for self-modifying agents, including reward hacking and self-termination (Orseau and Ring,, 2011). Open-ended empirical systems iterate code and toolchain edits with benchmark gains but without proof obligations (Zhang et al.,, 2025). Proposals for metagoals aim to stabilize or moderate goal evolution during self-change (Goertzel,, 2024). While these frameworks establish decision-theoretic foundations, they do not provide learning-theoretic guarantees about post-modification generalization. We study the learning-theoretic state of the agent after such edits.
2.2 Modern mechanisms for self-improvement
Contemporary machine learning exhibits constrained forms of self-modification across multiple dimensions. Neural architecture search explores architectural topologies through differentiable, evolutionary, and reinforcement approaches (Liu et al.,, 2019; Elsken et al.,, 2019; Zoph and Le,, 2017; Real et al.,, 2019). Automated machine learning systems perform pipeline and hyperparameter search and can trigger optimizer and model-family switches (Hutter et al.,, 2019; Feurer and Hutter,, 2019; Li et al.,, 2017). Population-based training simultaneously evolves hyperparameters and weights across a population of models (Jaderberg et al.,, 2017).
Meta-learning adapts optimizers, initializations, and inductive biases across tasks (Finn et al.,, 2017; Rajeswaran et al.,, 2019; Hospedales et al.,, 2022). Reinforcement learning and multi-armed bandits provide policies for selecting modifications and exploration strategies (Sutton and Barto,, 2018; Auer et al.,, 2002; Lai and Robbins,, 1985; Slivkins,, 2019). Representation growth through mixture of experts and adapters, and the use of external memory and retrieval, expand the effective function family and computation available at inference (Fedus et al.,, 2022; Houlsby et al.,, 2019; Hu et al.,, 2021; Graves et al.,, 2014, 2016; Lewis et al.,, 2020; Schick et al.,, 2023). Continual learning addresses sequential task acquisition while mitigating catastrophic forgetting (Kirkpatrick et al.,, 2017; Parisi et al.,, 2019; Van de Ven and Tolias,, 2022).
These mechanisms instantiate partial self-modification—adapting specific components while keeping the learning framework itself fixed. In contrast, true self-modifying agents can rewrite any axis of their design. In our framework, these mechanisms traverse representational, architectural, algorithmic, and metacognitive axes, and we establish when such traversals preserve or destroy learnability.
2.3 Learning theory for adaptive systems
PAC learning provides distribution-free guarantees under a fixed hypothesis class and algorithm (Shalev-Shwartz and Ben-David,, 2014; Mohri et al.,, 2018; Blumer et al.,, 1989; Vapnik,, 1998; Hanneke et al.,, 2024). Online learning theory establishes regret bounds for adaptive algorithms (Shalev-Shwartz,, 2012; Hazan,, 2016; Cesa-Bianchi and Lugosi,, 2006), but assumes the learning mechanism remains fixed. Transformation-invariant learners extend instance equivalence while keeping the learning mechanism fixed (Shao et al.,, 2022). Predictive PAC relaxes data assumptions with a fixed learner (Pestov,, 2010), and iterative improvement within constrained design spaces admits PAC-style analysis (Attias et al.,, 2025).
Information-theoretic approaches bound generalization for adaptive and meta-learners via mutual information (Jose and Simeone,, 2021; Chen et al.,, 2021; Wen et al.,, 2025). Stability connects optimization choices to generalization bounds (Bousquet and Elisseeff,, 2002; Hardt et al.,, 2016). All of these results assume architectural invariance: the hypothesis class, update rule, or computational model is fixed ex ante. We remove this assumption and characterize when self-modification preserves PAC learnability.
2.4 Computability and the substrate
Church–Turing-equivalent substrates preserve solvability up to simulation overhead, whereas strictly weaker substrates with finite memory can forfeit learnability of classes that are otherwise PAC-learnable; stronger-than-Turing models change the problem class under discussion (Akbari and Harrison-Trainor,, 2024). This motivates treating substrate edits separately from architectural or representational changes and clarifies when invariance should be expected.
3 Setup and Five-Axis Decomposition
At time $t\in\mathbb{N}$ the learner state is
$\ell_{t}=(A_{t},H_{t},Z_{t},F_{t},M_{t})\in\underbrace{\mathcal{L}}_{\mathcal{A}\times\mathcal{H}\times\mathcal{Z}\times\mathcal{F}\times\mathcal{M}},$
where $A$ (algorithmic), $H$ (representational), $Z$ (architectural), $F$ (substrate), and $M$ (metacognitive) are the five axes.
A possibly stochastic modification map $\Phi:\mathcal{L}\times\mathcal{D}\!\to\!\mathcal{L}$ updates the system via $\ell_{t+1}=\Phi(\ell_{t},D_{t})$. For $X\!\in\!\{A,H,Z,F,M\}$ with state space $\mathcal{X}$,
$X_{t+1}=\Phi_{X}(X_{t},D_{t},\theta_{X,t}),\qquad\theta_{X,t}\in\Theta_{X}.$
A reasonable utility $u$ is computable from finite state and data (for example empirical or validation risk with optional resource or complexity terms) and normalized to $[0,1]$. A state (and its components) is policy-reachable under $u$ if it can be obtained by iterating $\Phi$ from $\ell_{0}$ using the following decision rule.
Decision rule.
A candidate modification at time $t$ is executed iff there is a formal proof in the agent’s current calculus that it yields an immediate utility increase:
$u\!\big(\Phi(\ell_{t},D_{t}),\mathrm{Env}_{t}\big)\;>\;u\!\big(\ell_{t},\mathrm{Env}_{t}\big).$
We make no assumptions about proof-search efficiency; our results concern the post-modification learner.
Policy-reachable families (general form).
For any axis $X\in\{A,H,Z,F,M\}$, let $\mathcal{X}_{\mathrm{reach}}(u)$ be the set of $X$-states appearing in policy-reachable trajectories under $u$. In particular,
$\mathcal{X}_{\mathrm{reach}}(u)=\big\{\,$ $X^{\prime}:\ \exists\,t\text{ along a policy-reachable }$
$\text{trajectory under }u\text{ and }X_{t}=X^{\prime}\,\big\}.$
Capacity notion.
Losses are bounded in $[0,1]$. Our general statements hold for any uniform capacity notion that yields distribution-free uniform convergence (e.g., VC-subgraph or pseudodimension). For concreteness we instantiate to the 011 loss, where capacity reduces to the VC dimension VC()($\cdot$). (All bounds remain valid if VC is replaced by $\mathrm{Pdim}$ or VC-subgraph with the usual constants.)
Constants and notation.
[…]
Axis isolation and substrate scope.
Throughout, we analyze one axis at a time while holding others fixed. Under Church–Turing-equivalent substrates $F$, learnability refers to classical (Turing-based) PAC; non-CT cases are treated separately in the substrate section §8.
Data-path edits and i.i.d. integrity.
Our PAC statements assume $S$ and $V$ are i.i.d. from $\mathcal{D}$ and independent of each other. Permitted data-path operations are those that preserve i.i.d. draws (e.g., additional i.i.d. samples, balanced but label-independent subsampling, or predeclared splits). If selection depends on labels or on $V$, standard importance-weighting or covariate-shift corrections must be used; otherwise guarantees may fail.
Axes.
  1. 1.
    Algorithmic. Update rules, schedules, stopping, and internal randomness; the hypothesis family is fixed.
  2. 2.
    Representational. Changes to the hypothesis class or encoding, such as feature maps, basis expansions, or unions and refinements.
  3. 3.
    Architectural. Topology and information flow, including wiring, routing, depth or width, and memory addressing.
  4. 4.
    Substrate. Computational model and memory semantics, such as the machine model and memory capacity or discipline.
  5. 5.
    Metacognitive. A scheduler that selects and approves modifications on an enabled axis.
Why this decomposition matters.
The five-axis decomposition provides a conceptual toolkit for reasoning about arbitrary self-modification. Any concrete self-improving system can be analyzed by identifying which axes it modifies and tracing the induced changes to the hypothesis family. This has three immediate benefits: (i) it unifies seemingly disparate mechanisms under a common analytical framework; (ii) it isolates which modifications actually affect learnability (those that change the reachable hypothesis family) from those that affect only computational efficiency; and (iii) it enables modular safety certification—one can verify capacity bounds axis-by-axis rather than re-analyzing entire systems from scratch.
Concretely, these five axes already cover modern practice. Algorithmic edits appear in online hyperparameter and optimizer rewrites (Jaderberg et al.,, 2017). Representational expansions include mixture of experts and parameter-efficient adapters (Fedus et al.,, 2022; Houlsby et al.,, 2019; Hu et al.,, 2021). Architectural rewiring is explored by neural architecture search (Liu et al.,, 2019; Elsken et al.,, 2019). External memory, retrieval, and tool use expand effective computation while keeping computability assumptions unchanged (Graves et al.,, 2014, 2016; Lewis et al.,, 2020; Schick et al.,, 2023). Reflective and self-feedback agents instantiate metacognitive scheduling (Shinn et al.,, 2023; Madaan et al.,, 2023; Yao et al.,, 2023; Wang et al.,, 2023). As self-modifying capabilities become standard in deployed systems, this framework offers a principled way to predict, diagnose, and control when self-improvement preserves or destroys generalization guarantees.
Standing Assumptions and Scope.
  1. 1.
    (A1) Data $(x,y)$ i.i.d. from fixed $\mathcal{D}$; training $S\sim P^{m}$ and validation $V\sim P^{n_{v}}$ are independent.
  2. 2.
    (A2) Loss $\ell\in[0,1]$.
  3. 3.
    (A3) Capacity is any uniform-convergence notion (VC/pseudodim/VC-subgraph); we instantiate VC where convenient.
  4. 4.
    (A4) When a computable proxy $B$ is used, it upper-bounds capacity: $B(\cdot)\geq\mathrm{cap}(\cdot)$.
  5. 5.
    (A5) Substrate semantics. If the substrate $F$ is Church–Turing equivalent, solvability/learnability are measured in the classical (Turing) sense; non-CT substrates (e.g., finite memory or oracular/analog) may alter this and are treated separately in §8.
  6. 6.
    (A6) Axis isolation. We analyze one axis at a time while holding the others fixed; multi-axis edits are discussed later.
  7. 7.
    (A7) Compute scope. We study sample complexity (learnability), not runtime, unless limits are intrinsic to $F$.
4 Representational Self-Modification ($\mathcal{M}_{H}$)
Setting (fixed vs. modifiable).
We analyze representational edits while holding the algorithmic procedure $A$, architecture $Z$, substrate $F$, and metacognitive rule $M$ fixed. At time $t$ the learner has representation $H_{t}$ and a representational edit is
$H_{t+1}=\Phi_{H}(H_{t},D_{t},\theta_{t}).$
Data follow $(x,y)\sim\mathcal{D}$ i.i.d.; the training set $S\sim\mathcal{D}^{m}$ and validation set $V\sim\mathcal{D}^{n_{v}}$ are independent. Loss $\ell\in[0,1]$; risk $R(h)=\mathbb{E}[\ell(h(x),y)]$; empirical risks $\widehat{R}_{S}$, $\widehat{R}_{V}$. Capacity is measured by $\mathrm{VC}(\cdot)$. A utility $u$ is reasonable if it is computable, normalized to $[0,1]$, and (U1) non-decreasing in empirical fit on the active finite evidence, and (U2) strictly increasing in a computable capacity bonus $g(\mathrm{VC})$ with $g^{\prime}>0$. The decision rule executes an edit only when an immediate utility increase is formally provable. Proofs are deferred to Appendix 12.
Reference family.
[…]
Policy-reachable family.
For fixed $u$,
$\mathcal{H}_{\mathrm{reach}}(u)=\{\,$ $H^{\prime}:\ \exists t\ \text{along a policy-reachable}$
$\text{trajectory under }u\ \text{with }H_{t}=H^{\prime}\,\}.$
Unbounded representational power (URP).
$(\mathcal{H},\Phi_{H})$ has URP if for every $m\in\mathbb{N}$ there exist $H,\theta,D$ with
$\mathrm{VC}\!\big(\Phi_{H}(H,D,\theta)\big)\ \geq\ m.$
Local-URP: for every $H$ there exists an edit with $\mathrm{VC}(\Phi_{H}(H,\cdot,\cdot))\geq\mathrm{VC}(H)+1$ that can fit the current finite sample.
Policy-level learnability boundary
Under (A1–A7), distribution-free learnability is preserved under representational self-modification iff
$\sup_{H^{\prime}\in\mathcal{H}_{\mathrm{reach}}(u)}\mathrm{VC}(H^{\prime})<\infty.$
Sketch. ($\Leftarrow$) A uniform capacity cap gives uniform convergence on a fixed capped reference family for all steps; ERM/AERM yields the standard VC rate for the terminal predictor. ($\Rightarrow$) If capacities along a reachable subsequence diverge, VC lower bounds preclude any distribution-free sample complexity. (Full proof in Appendix 12.)
Two-Gate finite-sample safety
Given $S$ ($|S|=m$) and independent $V$ ($|V|=n_{v}$), a candidate edit producing $H_{\mathrm{new}}$ is accepted only if
(Validation) $\quad\widehat{R}_{V}(h_{\mathrm{new}})\leq\widehat{R}_{V}(h_{\mathrm{old}})-(2\varepsilon_{V}+\tau),$
(Capacity) $\quad h_{\mathrm{new}}\in\mathcal{G}_{K(m)}\ \ \text{with}\ \ \mathrm{VC}(\mathcal{G}_{K(m)})\leq K(m),$
where $\varepsilon_{V}$ is chosen so that, with probability $\geq 1-\delta_{V}$ over $V$,
$\sup_{h\in\mathcal{G}_{K(m)}}\big|R(h)-\widehat{R}_{V}(h)\big|\leq\ \varepsilon_{V}$
$\ (\varepsilon_{V}\asymp\sqrt{(K(m)+\log(1/\delta_{V}))/n_{v}}).$
Then with probability at least $1-\delta_{V}-\delta$ over draws of $V$ and $S$: (i) each accepted edit decreases true risk by at least $\tau$; and
$R(h_{T})\ \leq\ \inf_{h\in\mathcal{G}_{K(m)}}R(h)\ +\ \tilde{O}\!\Big(\sqrt{(K(m)+\log(1/\delta))/m}\Big).$
Validation reuse (fixed ex ante).
The same validation set $V$ may be reused adaptively across many edits provided the capped reference family $\mathcal{G}_{K(m)}$ is fixed before seeing $V$ and the gate thresholds $K(m),\varepsilon_{V},\tau$ do not depend on $V$. If any of these are tuned using $V$, a fresh split or reusable holdout is required (see Appendix 12).
Figure 1: Representational axis $\mathcal{M}_{H}$. TwoGate accepts a few early edits and then plateaus at a lower test loss, while destructive policies continue modifying the hypothesis class and exhibit worsening generalization as complexity increases.
Probability bookkeeping.
All oracle inequalities are stated on the intersection of two events: (E1) the uniform validation event on $\mathcal{G}_{K(m)}$ (probability $\geq 1-\delta_{V}$) and (E2) the training-side uniform convergence event (probability $\geq 1-\delta$). By a union bound, the final probability is $\geq 1-\delta_{V}-\delta$ and does not depend on the number of accepted edits, since the bound is uniform over the fixed capped family.
Remark.
Under URP, utilities that reward empirical fit and even a slight increase in capacity can drive $\mathrm{VC}$ unbounded and destroy distribution-free learnability; see Appendix 12.
5 Architectural Self-Modification ($\mathcal{M}_{Z}$)
Setting and reduction (fixed vs. modifiable).
We analyze architectural edits while holding the learning algorithm $A$, substrate $F$, and metacognitive rule $M$ fixed. An architecture $Z\in\mathcal{Z}$ induces a hypothesis class $H(Z)\subseteq\mathcal{Y}^{\mathcal{X}}$. At time $t$, an architectural edit produces
$Z_{t+1}\;=\;\Phi_{Z}(Z_{t},D_{t},\vartheta_{t}),$
$h_{S}(Z_{t+1})\in\underset{h\in H(Z_{t+1})}{\operatorname{argmin}}\widehat{R}_{S}(h)\ \$
$\hskip 14.22636pt\text{(ERM in the induced class)}.$
Fix a reasonable utility $u$. Let the policy-reachable architectures and induced classes be
$\mathcal{Z}_{\mathrm{reach}}(u)$ $=\big\{\,Z^{\prime}:\exists t\text{ on some proof-triggered}$
$\qquad\text{trajectory from }Z_{0}\text{ under }u\text{ with }Z_{t}=Z^{\prime}\,\big\},$
$\mathcal{H}^{Z}_{\mathrm{reach}}(u)$ $=\{\,H(Z):Z\in\mathcal{Z}_{\mathrm{reach}}(u)\,\}.$
Utility realism.
The boundary and Two-Gate guarantees depend only on the capacity of the reachable family, not on explicit capacity rewards: even if $u$ has no bonus term, any policy that permits capacity-increasing edits can cross the boundary unless a cap (e.g., via $K(m)$) is enforced. The stronger “bonus” variant is used only for the destruction theorems.
Every run of $\mathcal{M}_{Z}$ therefore corresponds to a run of $\mathcal{M}_{H}$ over the induced family $\mathcal{H}^{Z}_{\mathrm{reach}}(u)$.
Architectural $\to$ representational reduction
For any reasonable $u$, every proof-triggered trajectory $Z_{0}\!\to\!Z_{1}\!\to\!\cdots$ induces a representational trajectory $H(Z_{0})\!\to\!H(Z_{1})\!\to\!\cdots$ over the fixed reference family $\mathcal{G}$, with ERM/AERM inside each accepted $H(Z_{t})$. Consequently $\mathcal{H}^{Z}_{\mathrm{reach}}(u)=\{H(Z):Z\in\mathcal{Z}_{\mathrm{reach}}(u)\}\subseteq\mathcal{G}$. Proof sketch. The decision semantics and utility are unchanged by renaming states from $Z$ to the induced $H(Z)$.
Architectural boundary via induced policy-reachable family
For any reasonable $u$, distribution-free PAC learnability under architectural self-modification is preserved iff
$\sup_{Z\in\mathcal{Z}_{\mathrm{reach}}(u)}\mathrm{VC}\!\big(H(Z)\big)\ \leq\ K\;<\ \infty.$
Equivalently, preservation holds iff
$\sup_{H^{\prime}\in\mathcal{H}^{Z}_{\mathrm{reach}}(u)}\mathrm{VC}(H^{\prime})\leq K.$
Proof.
Immediate by reduction to §4: apply the sharp boundary theorem for representation (Thm. 4) to the induced set $\mathcal{H}^{Z}_{\mathrm{reach}}(u)$. ∎
Reference family and proxy-cap subfamily.
Fix a single parameterized super-family $\mathcal{G}\subseteq\mathcal{Y}^{\mathcal{X}}$ that contains every induced class: $H(Z)\subseteq\mathcal{G}$ for all $Z$. For $K\in\mathbb{N}$ define the proxy-cap subfamily
$\mathcal{G}^{\mathrm{proxy}}_{K}\ :=\ \{\,h\in\mathcal{G}:\ \exists Z\text{ with }B(Z)\leq K$
$\text{ and }h\in H(Z)$ $\,\},$
[…]
Two-gate safety for architecture
[…]
Validation reuse. We fix $\mathcal{G}^{\mathrm{proxy}}_{K(m)}$, $K(m)$, and thresholds ex ante (before seeing $V$). If any choice is tuned on $V$, we use a fresh split or a reusable-holdout mechanism; all theorems apply to the fixed family (Appendix 13.3).
Local architectural edits (Local-URPZ).
We say $\mathcal{M}_{Z}$ has Local-URPZ if for every $Z$ there exists a computable edit $(\vartheta,D)$ such that $\mathrm{VC}\!\big(H(\Phi_{Z}(Z,D,\vartheta))\big)\geq\mathrm{VC}(H(Z))+1$ and the new class $H(\Phi_{Z}(Z,D,\vartheta))$ can interpolate the current finite evidence (e.g., fit $S$).
Robust destruction under Local-URPZ
Assume Local-URPZ. For any reasonable utility $u$ (non-decreasing in empirical fit on the active evidence and strictly increasing in a computable capacity bonus $g(\mathrm{VC})$), there exist a distribution $\mathcal{D}$ and sample size $m$ such that the proof-trigger repeatedly accepts local architectural edits that increase capacity, the induced reachable set $\mathcal{H}^{Z}_{\mathrm{reach}}(u)$ has unbounded VC, and distribution-free PAC learnability fails. Sketch. Each local step strictly increases $u$ (fit non-worse + $g$ increases), so the proof-trigger fires. Iteration yields unbounded VC; apply the necessity direction of Thm. 5.
Proxy-cap sufficiency (computable architectural bounds).
Often one has a computable upper bound $B(Z)$ on $\mathrm{VC}(H(Z))$ (e.g., for ReLU networks $B(Z)\!=\!c\,W(Z)\log W(Z)$ with $W(Z)$ parameters).
Proxy-cap two-gate guarantee
If $B(Z)\geq\mathrm{VC}(H(Z))$ for all $Z$ and the capacity gate enforces $B(Z_{\mathrm{new}})\leq K(m)$, then Corollary 5 holds with $\mathcal{G}_{K(m)}=\{\,h\in\mathcal{G}:\ \text{$h$ realizable by some $Z$ with }B(Z)\leq K(m)\,\}$ and with the same rate (the VC bound is applied to $\mathcal{G}_{K(m)}$). Proof. Since $H(Z_{\mathrm{new}})\subseteq\mathcal{G}_{K(m)}$ and $\mathrm{VC}(\mathcal{G}_{K(m)})\leq K(m)$ by construction, the two-gate proof (Appendix 12) applies verbatim. ∎
6 Metacognitive Self-Modification ($\mathcal{M}_{M}$)
Setting (fixed vs. modifiable).
We analyze metacognitive scheduling/filters $M$ while holding representation $H$, architecture $Z$, algorithm $A$, and substrate $F$ fixed; $M$ selects which edits to evaluate and applies acceptance/rejection using finite evidence.
Metacognitive scheduler/filter.
A metacognitive rule $M$ can (i) choose which candidate edit to evaluate, (ii) when/how often to evaluate (scheduling), and (iii) randomize; it then accepts/rejects using only finite evidence (e.g., $S,V$, capacity proxy, edit cost). Let
$\mathcal{H}^{M,H}_{\mathrm{reach}}(u)\$$:=\ \Big\{\,H^{\prime}:\Pr\big[\exists t\ \text{on a proof-triggered}$
$\qquad\text{trajectory from }H_{0}\text{ under }u\text{ filtered}$
$\qquad\text{by }M\text{ with }H_{t}=H^{\prime}\big]>0\,\Big\}.$
This is the M-filtered policy-reachable family. (Randomized $M$ is allowed; all guarantees below hold almost surely over $M$’s internal randomness.)
Boundary under metacognitive modifications.
For any reasonable $u$ and metacognitive filter/scheduler $M$, distribution-free PAC learnability is preserved iff
$\sup_{H^{\prime}\in\mathcal{H}^{M,H}_{\mathrm{reach}}(u)}\mathrm{VC}(H^{\prime})\ \leq\ K\;<\ \infty.$
Proof sketch. Apply the representational sharp boundary to the $M$-filtered family $\mathcal{H}^{M,H}_{\mathrm{reach}}(u)$; necessity follows by the same destruction argument when the supremum is unbounded (Appendix 12).
Two-Gate as metacognition (safety + rate).
If $M$ accepts only when (i) $\widehat{R}_{V}(h_{t+1})\leq\widehat{R}_{V}(h_{t})-(2\varepsilon_{V}+\tau)$ and (ii) $\mathrm{VC}(H_{t+1})\leq K(m)$, then with probability $\geq 1-\delta_{V}-\delta$ each accepted edit reduces true risk by $\geq\tau$, and
$R(h_{T})\ \leq\ \inf_{h\in\mathcal{G}_{K(m)}}R(h)\ +\ \tilde{O}\!\Big(\sqrt{\tfrac{K(m)+\log(1/\delta)}{m}}\Big),$
by Thm. 4.
Restorative metacognition (turning destructive utilities safe).
Suppose the unfiltered reachable family satisfies $\sup_{H^{\prime}\in\mathcal{H}_{\mathrm{reach}}(u)}\mathrm{VC}(H^{\prime})=\infty$ (destructive). There exists a metacognitive rule $M$ (e.g., Two-Gate with any computable nondecreasing schedule $K(m)$) such that $\sup_{H^{\prime}\in\mathcal{H}^{M,H}_{\mathrm{reach}}(u)}\mathrm{VC}(H^{\prime})\leq K<\infty$, hence learnability is preserved.
Proof sketch. $M$ simply rejects any proposal that violates the validation margin or the cap; the filtered family is contained in $\mathcal{G}_{K(m)}$ (Appendix 12) and the boundary theorem applies.
Edit efficiency under metacognitive margins.
Under the Two-Gate rule with margin $\tau>0$, along any accepted trajectory
$\#\{\text{accepted edits}\}$ $\leq\frac{R(h_{0})-R^{*}}{\tau},$
$\text{where}\quad R^{*}$ $=\inf_{h\in\mathcal{G}_{K(m)}}R(h).$
Proof. Each accepted edit decreases true risk by at least $\tau$; telescope from $R(h_{0})$ to the floor $R^{*}$.
Scheduling and randomness.
$M$ may be deterministic or randomized and may schedule when candidate edits are proposed. Because our deviation bound is uniform over the capped family, the safety and rate guarantees hold for every realized trajectory filtered by $M$; scheduling affects efficiency, not the boundary itself.
7 Algorithmic Self-Modification ($\mathcal{M}_{A}$)
Setting (fixed vs. modifiable).
We analyze algorithmic self-modification while holding the hypothesis class $H$, architecture $Z$, substrate $F$, and metacognitive rule $M$ fixed. The learner may change its learning algorithm or schedule (optimizer, step sizes, noise level, stopping rule, etc.), producing an updated training procedure $A_{t+1}=\Phi_{A}(A_{t},D_{t},\vartheta_{t})$. Given a training set $S$ of size $m$, the output predictor is $\hat{h}=\mathsf{Alg}(A,S,H)\in H$.
Takeaways.
Algorithmic edits cannot cure infinite capacity; on finite capacity, ERM/AERM preserves PAC; and a simple stability meta-policy (cap the step-mass $\sum_{t}\eta_{t}$) controls the generalization gap during algorithmic self-modification.
Figure 2: $\mathcal{M}_{A}$ (Algorithmic axis). Generalization gap (test-train loss) vs. cumulative step–mass $M_{T}=\sum_{t}\eta_{t}$ with a fixed hypothesis class. TwoGate halts updates at a preset budget $B(m)$ and keeps the gap small; Destructive continues updating and exhibits a larger, persistent gap.
No algorithmic cure for infinite VC
[…]
Finite-VC is sufficient with ERM/AERM.
[…]
$R(\hat{h})\ \leq\ \inf_{h\in H}R(h)\ +\ \tilde{O}\!\Big(\sqrt{\tfrac{K+\log(1/\delta)}{m}}\Big).$
Stability meta-policy (step-mass cap).
Assume bounded, Lipschitz, $\beta$-smooth losses (formalized in Appendix 14), and that training examples are sampled uniformly from $S$ during updates. Let a self-modified training run on $H$ use (projected) SGD-like updates with step sizes $\{\eta_{t}\}_{t=1}^{T}$. Define the step-mass $M_{T}:=\sum_{t=1}^{T}\eta_{t}$.
When algorithm changes invalidate ERM assumptions.
If self-modification preserves ERM/AERM over the fixed $H$, Prop. 7 applies directly. If not, we control generalization via algorithmic stability: under standard smooth/Lipschitz conditions, the expected generalization gap scales as $O\!\big((\sum_{t}\eta_{t})/m\big)$; see Thm. 7.
Algorithmic stability via step-mass
Under the conditions above, there exists a constant $C>0$ (problem dependent, independent of $m$) such that
$\mathbb{E}\big[R(\hat{h})-\widehat{R}_{S}(\hat{h})\big]\ \leq\ \frac{C}{m}\,\sum_{t=1}^{T}\eta_{t}\;=\;\frac{C}{m}\,M_{T}.$
In particular, a metacognitive rule that caps $M_{T}\leq B(m)$ guarantees $\mathbb{E}[\mathrm{gap}]=\tilde{O}(B(m)/m)$; choosing $B(m)=\tilde{O}(1)$ yields a $\tilde{O}(1/m)$ gap.
Discussion.
[…]
8 Substrate Self-Modification ($\mathcal{M}_{F}$)
Setting (fixed vs. modifiable).
We analyze substrate edits while holding the specification of $H$, $Z$, and $A$ fixed; switching substrates changes how these are executed (time/space) but not which hypotheses are definable nor which utilities are expressible.
Setting.
The learner runs on a computational substrate $F$ (hardware/VM). A substrate edit is
$F_{t+1}\;=\;\Phi_{F}(F_{t},D_{t},\varphi_{t}).$
Unless otherwise stated, the representation $H$, architecture $Z$, and algorithmic procedure $A$ are held fixed as specifications; switching substrates may change how they are executed (time/space), but not which hypotheses are definable nor which utilities are expressible. PAC learnability is classical (i.i.d., 011 loss unless noted).
Takeaways.
(i) Switching among Church–Turing equivalent substrates preserves classical PAC learnability (CT-invariance). (ii) Downgrading to a strictly weaker substrate (e.g., finite-state) can destroy PAC learnability even for problems learnable pre-switch. (iii) “Beyond-CT” substrates do not alter classical PAC guarantees unless they enlarge the induced hypothesis family; if they do, our policy-reachable boundary from §4 applies to the enlarged family.
CT-invariance of PAC learnability
If $F$ and $F^{\prime}$ are Church–Turing equivalent, then a problem that is distribution-freely PAC-learnable when run on $F$ remains distribution-freely PAC-learnable when run on $F^{\prime}$, with the same sample complexity up to constant factors (computation may differ).
Finite-state downgrade can destroy learnability
There exists a binary classification problem that is PAC-learnable on a Turing-equivalent substrate but becomes not distribution-freely PAC-learnable after switching to a fixed finite-state substrate (bounded persistent memory), even when $H$ has finite VC dimension as a specification.
Beyond-CT substrates
If a stronger-than-Turing substrate $F^{\dagger}$ does not enlarge the induced hypothesis family (the measurable set of predictors under consideration), classical PAC learnability is unchanged. If it does enlarge the effective family to $H^{\dagger}$, the learnability boundary is governed by $\sup_{H^{\prime}\in\mathcal{H}^{\dagger}_{\mathrm{reach}}(u)}\text{VC}(H^{\prime})$ exactly as in §4.
Discussion.
Thm. 8 elevates “substrate choice” out of the classical PAC calculus: CT-equivalent machines only affect compute, not sample complexity. Prop. 8 formalizes the “simpleton downgrading” intuition: collapsing persistent memory imposes an information bottleneck that breaks distribution-free guarantees. Prop. 8 shows that any real change to learnability arises only through the induced hypothesis family; our policy-reachable boundary applies verbatim once that family changes. Full proofs appear in Appendix 15.
9 Outlook
From theory to practice: Why capacity bounds matter.
Practitioners may object that modern deep learning routinely violates PAC sample complexity through implicit regularization—why would capacity bounds matter for self-modifying systems? The answer is sequential compounding risk. A single overparameterized model may generalize well, but a self-modifying agent that repeatedly expands capacity across hundreds of edits accumulates risk that no implicit bias can rescue. Our experiments demonstrate this concretely: the destructive policy reaches test loss 0.409 after unbounded capacity growth, while the capacity-capped policy achieves 0.350—a 17% relative improvement emerging not from a single edit, but from accumulated drift as the reachable family explodes. The Two-Gate policy translates directly into practice: (i) track a capacity proxy $B(\cdot)$ (parameter count for neural architectures, corpus size for retrieval systems, function-library size for tool-using agents); (ii) set a schedule $K(m)\propto\sqrt{m}$ proportional to available training data; (iii) reject edits unless validation improves by margin $\tau$. These checks are computationally cheap and can run in real-time during self-improvement loops. The alternative to capacity bounds is not ”trust implicit regularization”—it is accepting that your system has entered a regime where no learning guarantee is possible.
Multi-axis modification: The realistic frontier.
Real autonomous agents will simultaneously rewrite architectures, swap optimizers, expand tool libraries, and adjust metacognitive policies—not modify a single axis in isolation. Our framework extends naturally: all axes reduce to the same boundary condition on $\sup\mathrm{VC}(\mathcal{H}_{\mathrm{reach}})$. When an agent modifies multiple axes, the induced hypothesis family is determined by the composition of edits—e.g., architecture $Z^{\prime}$ combined with representation $H^{\prime}$ induces $H^{\prime}(Z^{\prime})$, and learnability requires the joint trajectory remain capacity-bounded. This has three practical implications. First, capacity bounds must be enforced globally, not per-axis—a small architectural change combined with representational expansion can yield multiplicative VC growth. Second, axis interactions create emergent capacity explosions that independent per-axis budgets cannot prevent. Third, metacognitive policies become essential: in multi-axis settings, compounding accelerates capacity growth exponentially, making global capacity monitoring the only known mechanism for ex ante safety. The key open challenge is developing compositional capacity proxies that tractably upper-bound $\mathrm{VC}(H^{\prime}(Z^{\prime})\circ\text{tools})$ for complex compositions—the gap between computable bounds $B(\cdot)$ and true VC dimension determines how conservative the Two-Gate policy must be.
Towards sustainable self-improvement.
The AutoML and NAS communities have achieved remarkable success by treating architecture search as unconstrained optimization. Our results suggest a paradigm shift: future self-modifying systems should ask not ”what maximizes validation accuracy?” but ”what maximizes accuracy subject to capacity remaining PAC-learnable for available data?” This constraint does not eliminate innovation—it channels self-improvement toward sustainable compounding of gains rather than compounding of risk. For open-ended agents operating over months or years, the capacity schedule $K(m)$ can grow with accumulating data $m$, enabling unbounded absolute improvement while maintaining learnability. The failure mode is not self-improvement itself, but uncontrolled self-improvement that outruns the data. Without capacity bounds, seemingly rational modifications can lock in poor performance irreversibly—once an agent crosses into the unbounded regime, no sample complexity guarantees recovery. For high-stakes deployments (medical AI, autonomous vehicles, financial systems), the choice is between principled capacity-aware self-modification and accepting that you have no basis for trust.
10 Conclusion
We have established a sharp learnability boundary for self-modifying agents: distribution-free PAC guarantees are preserved if and only if the policy-reachable hypothesis family has uniformly bounded capacity. This result unifies representational, architectural, algorithmic, metacognitive, and substrate modifications under a single criterion—the supremum VC dimension of states reachable under the agent’s utility function. The Two-Gate policy provides a computable guardrail that enforces this boundary through validation margins and capacity caps, yielding oracle inequalities at standard VC rates. Our framework reveals that seemingly rational self-modifications can irreversibly destroy learnability when capacity grows without bound, even as they improve immediate performance. As AI systems gain the capability to rewrite their own learning mechanisms, the choice is between principled capacity-aware self-improvement that preserves generalization guarantees, or unconstrained optimization that enters a regime where no learning theory can provide safety assurances. The path to safe open-ended autonomy requires recognizing that self-modification is not a bug to eliminate, but a capability to control—and capacity bounds are the control mechanism that learning theory provides.
Appendix
11 Definitions
Symbol Definitions
$H(Z)$ Hypothesis class induced by architecture $Z$
$\mathcal{H}_{\mathrm{reach}}(u)$ Policy-reachable hypothesis family under utility $u$
$B(\cdot)$ Computable capacity proxy (upper-bounds VC/pseudodim)
$K(m)$ Nondecreasing capacity cap schedule at sample size $m$
$\mathcal{G}_{K(m)}$ Reference subfamily with $\mathrm{cap}\leq K(m)$
$\tau$ Validation margin in Two-Gate
$\widehat{R}_{S},\widehat{R}_{V}$ Empirical risks on train $S$ and validation $V$
$R^{*}$ $\inf_{h\in\mathcal{G}_{K(m)}}R(h)$
Table 1: Notation.
12 Full Proofs for Representational Self-Modification
Data, loss, risks.
Samples $(x,y)\sim\mathcal{D}$ i.i.d. Training $S\sim\mathcal{D}^{m}$ and validation $V\sim\mathcal{D}^{n_{v}}$ are independent. Loss $\ell\in[0,1]$; true risk $R(h)=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(h(x),y)]$. Empirical risks are $\widehat{R}_{S}$ and $\widehat{R}_{V}$ on $S$ and $V$.
Representational edits and policies.
At time $t$ the representation is a hypothesis class $H_{t}\subseteq\mathcal{Y}^{\mathcal{X}}$. A representational edit is $H_{t+1}=\Phi_{H}(H_{t},D_{t},\theta_{t})$, where $D_{t}$ is finite evidence (e.g., $S$, summary stats) and $\theta_{t}$ are edit parameters. Within any accepted class $H$, the learner outputs an ERM (or AERM) on $S$:
$h_{S}(H)\in\arg\min_{h\in H}\widehat{R}_{S}(h).$
The decision rule executes an edit only when an immediate utility increase is formally provable from the finite evidence.
Reasonable utilities.
A utility $u$ is reasonable if it is (i) computable from finite state/evidence and (ii) satisfies:
  • (U1) Non-decreasing in empirical fit on the active finite evidence (e.g., $1-\widehat{R}_{S}$).
  • (U2) Adds a strictly increasing capacity bonus $g(\mathrm{VC}(H))$ with $g^{\prime}(k)>0$ for all $k$.
We normalize $u\in[0,1]$ WLOG.
Policy-reachable family.
Fix $u$. Let $\mathcal{H}_{\mathrm{reach}}(u)$ be the set of classes $H^{\prime}$ for which there exists a time $t$ on some proof-triggered trajectory from $H_{0}$ under $u$ with $H_{t}=H^{\prime}$.
URP and Local-URP.
The pair $(\mathcal{H},\Phi_{H})$ has URP if for every $m\in\mathbb{N}$ there exist $(H,\theta,D)$ with $\mathrm{VC}(\Phi_{H}(H,D,\theta))\geq m$. It has Local-URP if for every $H$ there is an edit $(\theta,D)$ such that $\mathrm{VC}(\Phi_{H}(H,D,\theta))\geq\mathrm{VC}(H)+1$ and the new class can interpolate the current finite evidence (e.g., fit $S$).
Single capped reference family (indexing control).
For each $K\in\mathbb{N}$ let $\mathcal{G}_{K}\subseteq\mathcal{Y}^{\mathcal{X}}$ be a reference family with $\mathrm{VC}(\mathcal{G}_{K})\leq K$ and assume the capacity gate (defined below) guarantees all accepted $H$ satisfy $H\subseteq\mathcal{G}_{K}$. This avoids pathologies from taking unions over arbitrarily many distinct classes with the same VC cap.
VC uniform convergence.
There exists a universal constant $c>0$ such that for any class $G$ with $\mathrm{VC}(G)\leq K$ and any $\delta\in(0,1)$, with probability $\geq 1-\delta$ over a sample of size $n$,
$\sup_{h\in G}\big|R(h)-\widehat{R}(h)\big|\ \leq\ c\sqrt{\frac{K+\log(1/\delta)}{n}}.$ (1)
[…]
12.1 Sharp policy-level boundary
Theorem (Sharp boundary; restated from Thm. 4).
For any reasonable $u$, distribution-free PAC learnability is preserved under representational self-modification iff there exists $K<\infty$ such that
$\sup_{H^{\prime}\in\mathcal{H}_{\mathrm{reach}}(u)}\mathrm{VC}(H^{\prime})\ \leq\ K.$
Proof (sufficiency).
Fix $u$ and assume $\sup_{H^{\prime}\in\mathcal{H}_{\mathrm{reach}}(u)}\mathrm{VC}(H^{\prime})\leq K$. Along any policy-reachable run, all classes satisfy $H_{t}\subseteq\mathcal{G}_{K}$ with $\mathrm{VC}(\mathcal{G}_{K})\leq K$. By (1) applied to $\mathcal{G}_{K}$ and ERM in $H_{t}$,
$R\big(h_{S}(H_{t})\big)\ \leq\ \inf_{h\in H_{t}}R(h)\ +\ \tilde{O}\!\Big(\sqrt{\tfrac{K+\log(1/\delta)}{m}}\Big)$
uniformly for all $t$, with probability $\geq 1-\delta$ over $S$. In particular the terminal predictor $h_{T}$ obeys the same bound, so $m=\tilde{O}\!\big((K+\log(1/\delta))/\epsilon^{2}\big)$ suffices for $(\epsilon,\delta)$-accuracy. ∎
Proof (necessity).
If $\sup_{H^{\prime}\in\mathcal{H}_{\mathrm{reach}}(u)}\mathrm{VC}(H^{\prime})=\infty$, then for each $k$ there exists a reachable $H^{(k)}$ with $\mathrm{VC}(H^{(k)})\geq k$. Classical VC lower bounds imply any distribution-free learner needs $m=\Omega(k/\epsilon)$ samples for $(\epsilon,\delta)$-accuracy (even realizable). Since $k$ is unbounded along reachable trajectories, no uniform PAC guarantee exists. ∎
12.2 Finite-sample safety of the two-gate policy
Two gates.
Given train $S$ ($|S|=m$) and independent validation $V$ ($|V|=n_{v}$), accept an edit only if:
$\textbf{(Validation)}\quad\widehat{R}_{V}(h_{\mathrm{new}})\ \leq\ \widehat{R}_{V}(h_{\mathrm{old}})\ -\ (2\varepsilon_{V}+\tau),$
$\textbf{(Capacity)}\quad H_{\mathrm{new}}\ \subseteq\ \mathcal{G}_{K(m)}\qquad\text{with}\quad\mathrm{VC}(\mathcal{G}_{K(m)})\leq K(m),$
where $K(\cdot)$ is nondecreasing, $\tau\geq 0$ is a margin, and $\varepsilon_{V}$ is chosen so that with probability $\geq 1-\delta_{V}$ over $V$,
[…]
Theorem (Two-gate finite-sample safety; restated from Thm. 4).
With probability $\geq 1-\delta_{V}-\delta$ over $(V,S)$:
(i) each accepted edit decreases true risk by at least $\tau$ (monotone steps); and
(ii) the terminal predictor $h_{T}$ satisfies the oracle inequality
$R(h_{T})\ \leq\ \inf_{h\in\mathcal{G}_{K(m)}}R(h)\ +\ \tilde{O}\!\Big(\sqrt{\tfrac{K(m)+\log(1/\delta)}{m}}\Big).$
Proof.
(i) On the event $\sup_{h\in\mathcal{G}_{K(m)}}|R(h)-\widehat{R}_{V}(h)|\leq\varepsilon_{V}$,
$R(h_{\mathrm{new}})\ \leq\ \widehat{R}_{V}(h_{\mathrm{new}})+\varepsilon_{V}\ \leq\ \widehat{R}_{V}(h_{\mathrm{old}})-(2\varepsilon_{V}+\tau)+\varepsilon_{V}\ \leq\ R(h_{\mathrm{old}})-\tau.$
(ii) By the capacity gate, $h_{T}\in\mathcal{G}_{K(m)}$. Apply (1) to $\mathcal{G}_{K(m)}$ on $S$ and ERM in $H_{T}$ to obtain
$R(h_{T})\ \leq\ \inf_{h\in\mathcal{G}_{K(m)}}R(h)\ +\ \tilde{O}\!\Big(\sqrt{\tfrac{K(m)+\log(1/\delta)}{m}}\Big)$
with probability $\geq 1-\delta$. Union bound with the validation event yields the stated probability. ∎
12.3 Destruction under (local) URP
Theorem (Existential destruction under URP).
Assume URP. There exists a reasonable utility $u$ and a problem that is PAC-learnable in the baseline class such that the proof-triggered policy executes representational edits that render the problem distribution-free unlearnable after modification.
Proof.
Let $C$ be any finite-VC concept class (PAC-learnable without modification). Define a computable utility $u=\alpha\,(1-\widehat{R}_{S}(h))+\beta\,g(\mathrm{VC}(H))$ with $\alpha,\beta>0$ and strictly increasing $g$. By URP, for the realized $S$ there exists an edit to $H^{\star}$ with $\mathrm{VC}(H^{\star})\geq|S|$ that interpolates $S$ (shatters $S$). Then $u$ strictly increases (perfect fit plus larger capacity bonus), which is provable from finite evidence; the proof-trigger executes the edit. With $\mathrm{VC}(H^{\star})\geq|S|$ and only $|S|$ samples, standard VC lower bounds show distribution-free PAC learnability fails. ∎
Theorem (Utility-class robust destruction under Local-URP).
Assume Local-URP. Then for any reasonable utility $u$ (satisfying (U1)–(U2)), there exist a distribution $\mathcal{D}$ and sample size $m$ such that the proof-trigger repeatedly accepts capacity-increasing local edits, the policy-reachable VC is unbounded, and distribution-free PAC learnability fails.
Proof.
Local-URP ensures at each step a computable edit with $\mathrm{VC}$ increase by at least $1$ that preserves or improves empirical fit on the active evidence. By (U1)–(U2), the utility strictly increases at each such step, hence the proof-trigger fires. Iteration yields unbounded policy-reachable VC, so by the necessity part of the boundary theorem learnability cannot be guaranteed distribution-free. ∎
13 Full Proofs for Architectural Self-Modification
Assumptions & tools (recap).
Data $(x,y)\sim\mathcal{D}$ i.i.d.; training $S\sim\mathcal{D}^{m}$ and validation $V\sim\mathcal{D}^{n_{v}}$ are independent. Loss $\ell\in[0,1]$; true risk $R(h)=\mathbb{E}[\ell(h(x),y)]$; empirical risks $\widehat{R}_{S},\widehat{R}_{V}$ on $S,V$. Within any accepted class $H$, the learner outputs ERM $h_{S}(H)\in\arg\min_{h\in H}\widehat{R}_{S}(h)$. A utility $u$ is reasonable if it is computable from finite evidence and satisfies: (U1) non-decreasing in empirical fit on the active evidence (e.g., $1-\widehat{R}_{S}$), and (U2) strictly increasing in a computable capacity bonus $g(\mathrm{VC}(H))$ with $g^{\prime}(k)>0$. We use the standard VC uniform-convergence bound: for a class $G$ with $\mathrm{VC}(G)\leq K$ and any $\delta\in(0,1)$,
$\sup_{h\in G}\big|R(h)-\widehat{R}(h)\big|\ \leq\ c\sqrt{\tfrac{K+\log(1/\delta)}{n}}\qquad(\text{with a universal constant }c>0).$ (2)
[…]
Architectures induce classes and a reference family.
An architecture $Z\in\mathcal{Z}$ induces a hypothesis class $H(Z)\subseteq\mathcal{Y}^{\mathcal{X}}$. We assume a single reference family $\mathcal{G}\subseteq\mathcal{Y}^{\mathcal{X}}$ such that $H(Z)\subseteq\mathcal{G}$ for all $Z$. For each $K\in\mathbb{N}$, let $\mathcal{G}_{K}\subseteq\mathcal{G}$ be a subfamily with $\mathrm{VC}(\mathcal{G}_{K})\leq K$. (Concrete choices include $\mathcal{G}$ the realizable functions of a supernet, and $K$ derived from an architectural proxy such as parameter count; see Prop. 13.3.)
Policy reachability in architecture.
Fix a reasonable utility $u$ and the proof-triggered decision rule. Define
[…]
and the induced family
$\mathcal{H}^{Z}_{\mathrm{reach}}(u)\ =\ \big\{\,H(Z):\ Z\in\mathcal{Z}_{\mathrm{reach}}(u)\,\big\}\ \subseteq\ \mathcal{G}.$
13.1 Reduction lemma: $\mathcal{M}_{Z}\rightarrow\mathcal{M}_{H}$
Lemma (Architectural-to-representational reduction).
Fix a reasonable utility $u$. Any proof-triggered trajectory $Z_{0}\to Z_{1}\to\cdots$ in $\mathcal{M}_{Z}$ induces a trajectory $H(Z_{0})\to H(Z_{1})\to\cdots$ in $\mathcal{M}_{H}$ over the family $\mathcal{H}^{Z}_{\mathrm{reach}}(u)$, with predictors $h_{S}(Z_{t})\in\arg\min_{h\in H(Z_{t})}\widehat{R}_{S}(h)$ (ERM).
Proof.
By definition, at time $t$ the available hypotheses are exactly $H(Z_{t})$. The learner outputs ERM within $H(Z_{t})$. The proof-triggered rule accepts an edit iff there exists a formal proof (from finite evidence) that $u$ increases. Because $u$ and the decision semantics are identical whether we name the state by $Z_{t}$ or by $H(Z_{t})$, each accepted architectural edit corresponds to an accepted representational edit on the induced class, and vice versa. Thus the reachable set under $u$ maps to $\mathcal{H}^{Z}_{\mathrm{reach}}(u)$, and the architectural run induces a representational run along the mapped classes.
Corollary (Architectural boundary by reduction).
For any reasonable $u$, distribution-free PAC learnability under architectural self-modification is preserved iff $\sup_{Z\in\mathcal{Z}_{\mathrm{reach}}(u)}\mathrm{VC}(H(Z))\leq K<\infty$.
Proof.
Apply the sharp boundary theorem for representation (Appendix 12.1) to the induced set $\mathcal{H}^{Z}_{\mathrm{reach}}(u)$ using Lemma 13.1. See also the main-text statement Thm. 5.
13.2 Local architectural edits and robust destruction
Local-URPZ.
We say $\mathcal{M}_{Z}$ has Local-URPZ if for every $Z$ there exists a computable edit $(\vartheta,D)$ such that: (i) $\mathrm{VC}\!\big(H(\Phi_{Z}(Z,D,\vartheta))\big)\ \geq\ \mathrm{VC}(H(Z))+1$; and (ii) the new class $H(\Phi_{Z}(Z,D,\vartheta))$ can interpolate the active finite evidence (e.g., fit $S$).
Theorem (Robust destruction under Local-URPZ).
Assume Local-URPZ. For any reasonable utility $u$ satisfying (U1)–(U2), there exist a distribution $\mathcal{D}$ and sample size $m$ such that the proof-trigger repeatedly accepts local architectural edits that increase capacity, the induced reachable set $\mathcal{H}^{Z}_{\mathrm{reach}}(u)$ has unbounded VC, and distribution-free PAC learnability fails.
Proof.
Fix $u$. From any $Z_{t}$, Local-URPZ guarantees a computable edit $(\vartheta_{t},D_{t})$ producing $Z_{t+1}$ with $\mathrm{VC}(H(Z_{t+1}))\geq\mathrm{VC}(H(Z_{t}))+1$ and perfect fit to the active finite evidence (e.g., $S$). By (U1) the empirical fit term in $u$ is non-worse; by (U2) the capacity bonus $g(\mathrm{VC})$ strictly increases; the net increase in $u$ is a computable fact from finite evidence. Therefore the proof-trigger fires and the edit is accepted. Iterating yields a reachable sequence with unbounded $\mathrm{VC}(H(Z_{t}))$. Hence $\sup_{H^{\prime}\in\mathcal{H}^{Z}_{\mathrm{reach}}(u)}\mathrm{VC}(H^{\prime})=\infty$, and by the necessity direction of the sharp boundary (Appendix 12.1; cf. Thm. 4) distribution-free PAC learnability cannot be guaranteed.
13.3 Proxy-capacity two-gate guarantee
Computable architectural upper bounds.
Suppose there exists a computable function $B:\mathcal{Z}\to\mathbb{N}$ with
$\mathrm{VC}\!\big(H(Z)\big)\ \leq\ B(Z)\qquad\text{for all }Z\in\mathcal{Z}.$ (3)
(Examples: for ReLU networks, $B(Z)\!=\!c\,W(Z)\log W(Z)$ with $W(Z)$ the parameter count.)
Proxy-capped reference subfamily.
For $K\in\mathbb{N}$ define
$\mathcal{G}^{\mathrm{proxy}}_{K}\ :=\ \Big\{\,h\in\mathcal{G}:\ \exists Z\in\mathcal{Z}\text{ with }B(Z)\leq K\text{ and }h\in H(Z)\,\Big\}.$
By (3), $\mathrm{VC}\!\big(\mathcal{G}^{\mathrm{proxy}}_{K}\big)\leq K$.
Proposition (Proxy-cap two-gate oracle inequality).
Assume the two-gate policy with capacity gate $B(Z_{\mathrm{new}})\leq K(m)$ and validation gate
$\widehat{R}_{V}\!\big(h_{S}(Z_{\mathrm{new}})\big)\ \leq\ \widehat{R}_{V}\!\big(h_{S}(Z_{\mathrm{old}})\big)\ -\ (2\varepsilon_{V}+\tau),$
where $\varepsilon_{V}$ is chosen so that $\sup_{h\in\mathcal{G}^{\mathrm{proxy}}_{K(m)}}|R(h)-\widehat{R}_{V}(h)|\leq\varepsilon_{V}$ with probability $\geq 1-\delta_{V}$ (e.g., by (2)). Then with probability $\geq 1-\delta_{V}-\delta$ over $(V,S)$:
  1. 1.
    (Monotone steps) Each accepted architectural edit satisfies $R\!\big(h_{S}(Z_{\mathrm{new}})\big)\ \leq\ R\!\big(h_{S}(Z_{\mathrm{old}})\big)-\tau$.
  2. 2.
    (Oracle inequality) The terminal predictor $h_{S}(Z_{T})$ obeys
    $R\!\big(h_{S}(Z_{T})\big)\ \leq\ \inf_{h\in\mathcal{G}^{\mathrm{proxy}}_{K(m)}}R(h)\ +\ \tilde{O}\!\Big(\sqrt{\tfrac{K(m)+\log(1/\delta)}{m}}\Big).$
Proof.
By the capacity gate and the definition of $\mathcal{G}^{\mathrm{proxy}}_{K(m)}$, every accepted class $H(Z_{\mathrm{new}})$ is a subset of $\mathcal{G}^{\mathrm{proxy}}_{K(m)}$ with VC bounded by $K(m)$. The monotone-step claim follows exactly as in the representational two-gate proof: on the high-probability validation event,
$R(h_{\mathrm{new}})\ \leq\ \widehat{R}_{V}(h_{\mathrm{new}})+\varepsilon_{V}\ \leq\ \widehat{R}_{V}(h_{\mathrm{old}})-(2\varepsilon_{V}+\tau)+\varepsilon_{V}\ \leq\ R(h_{\mathrm{old}})-\tau.$
For the oracle inequality, apply (2) on $S$ to the capped family $\mathcal{G}^{\mathrm{proxy}}_{K(m)}$ and use ERM in the final accepted class. Union bound the training and validation events to obtain probability $\geq 1-\delta_{V}-\delta$.
Remark (using a fixed reference subfamily).
If the capacity gate is stated directly as $H(Z_{\mathrm{new}})\subseteq\mathcal{G}_{K(m)}$ with $\mathrm{VC}(\mathcal{G}_{K(m)})\leq K(m)$, then Proposition 13.3 holds verbatim with $\mathcal{G}^{\mathrm{proxy}}_{K(m)}$ replaced by $\mathcal{G}_{K(m)}$.
13.4 Pointers back to representational results
Lemma 13.1 allows Theorem 5 in the main text to be proved by direct appeal to the representational sharp boundary (Appendix 12.1; cf. Thm. 4). Corollary 5 (main text) is an instantiation of the two-gate safety theorem (Appendix 12.2) with the architectural capacity gate $H(Z)\subseteq\mathcal{G}_{K(m)}$.
14 Full Proofs for Algorithmic Self-Modification
Assumptions.
Data $(x,y)\sim\mathcal{D}$ i.i.d.; $S\sim\mathcal{D}^{m}$. Loss $\ell(\cdot;\,z)$ is bounded in $[0,1]$, $L$-Lipschitz in the parameter $\theta$ (with respect to a norm $\|\cdot\|$), and $\beta$-smooth. Gradients are bounded $\|\nabla_{\theta}\ell(\theta;z)\|\leq G$, or we use projection onto a bounded domain of diameter $D$ so iterates remain bounded. The hypothesis class $H=\{x\mapsto f_{\theta}(x):\ \theta\in\Theta\}$ is fixed throughout the algorithmic edits. When we invoke ERM/AERM, $\hat{h}\in\arg\min_{h\in H}\widehat{R}_{S}(h)$ (or an approximate minimizer).
14.1 No algorithmic cure for infinite VC
[…]
14.2 ERM/AERM on finite VC
[…]
$\sup_{h\in H}\big|R(h)-\widehat{R}_{S}(h)\big|\ \leq\ c\sqrt{\tfrac{K+\log(1/\delta)}{m}}$
for a universal constant $c>0$ (polylogs hidden in $\tilde{O}$). If $\hat{h}$ is ERM/AERM in $H$, then
$R(\hat{h})\ \leq\ \widehat{R}_{S}(\hat{h})+\tilde{O}\!\Big(\sqrt{\tfrac{K+\log(1/\delta)}{m}}\Big)\ \leq\ \inf_{h\in H}\widehat{R}_{S}(h)+\tilde{O}\!\Big(\sqrt{\tfrac{K+\log(1/\delta)}{m}}\Big)$
$\leq\ \inf_{h\in H}R(h)\ +\ \tilde{O}\!\Big(\sqrt{\tfrac{K+\log(1/\delta)}{m}}\Big).$
Thus ERM/AERM preserves the PAC rate on a fixed finite-VC class. ∎
14.3 Proof of Thm. 7: stability via step-mass
We prove a uniform stability bound for (projected) SGD-like updates and then translate it to an expected generalization bound.
Algorithm and neighboring samples.
Let $S=(z_{1},\dots,z_{m})$ and $S^{(i)}$ be $S$ with the $i$th example replaced by an independent copy $z^{\prime}_{i}$. Run the same (possibly self-modified) algorithmic schedule on $S$ and $S^{(i)}$ with shared randomness. Denote parameter sequences $\{\theta_{t}\}_{t=0}^{T}$ and $\{\theta^{\prime}_{t}\}_{t=0}^{T}$ with updates
$\theta_{t+1}\ =\ \Pi\!\left(\theta_{t}-\eta_{t}g_{t}\right),\qquad g_{t}\in\partial\ell(\theta_{t};z_{I_{t}}),$
where $\Pi$ is projection (if used) and $I_{t}$ is the sampled index at step $t$ (uniform on $[m]$ or any scheme that samples from $S$); define $\theta^{\prime}_{t+1}$ analogously with $S^{(i)}$ and the same $I_{t}$ sequence.
One-step sensitivity.
By nonexpansiveness of projection and $\beta$-smoothness with step sizes $\eta_{t}\leq 1/\beta$,
$\|\theta_{t+1}-\theta^{\prime}_{t+1}\|\ \leq\ \|\theta_{t}-\theta^{\prime}_{t}\|+\eta_{t}\|g_{t}-g^{\prime}_{t}\|.$
If $I_{t}\neq i$ then $z_{I_{t}}$ is identical in both runs and $\|g_{t}-g^{\prime}_{t}\|\leq L\|\theta_{t}-\theta^{\prime}_{t}\|$ (by $L$-Lipschitzness of gradients under smoothness). If $I_{t}=i$, the gradients can differ by at most $2G$. Taking conditional expectation over $I_{t}$ (uniform sampling) yields
$\mathbb{E}\big[\|\theta_{t+1}-\theta^{\prime}_{t+1}\|\ \big|\ \theta_{t},\theta^{\prime}_{t}\big]\ \leq\ \Big(1+\tfrac{L}{m}\eta_{t}\Big)\|\theta_{t}-\theta^{\prime}_{t}\|\ +\ \tfrac{2G}{m}\eta_{t}.$
Iterating from identical initialization gives
$\mathbb{E}\|\theta_{T}-\theta^{\prime}_{T}\|\ \leq\ \tfrac{2G}{m}\sum_{t=1}^{T}\eta_{t}\prod_{s=t+1}^{T}\Big(1+\tfrac{L}{m}\eta_{s}\Big)\ \leq\ \tfrac{2G}{m}\,e^{(L/m)\sum_{s}\eta_{s}}\sum_{t=1}^{T}\eta_{t}\ \leq\ \tfrac{2Ge^{L}}{m}\sum_{t=1}^{T}\eta_{t},$
where we used $\sum_{s}\eta_{s}\leq M_{T}$ and $M_{T}/m\leq 1$ for the exponent (or simply absorb $e^{(L/m)M_{T}}$ into the constant).
From parameter sensitivity to loss stability.
By $L$-Lipschitzness of $\ell(\cdot;z)$ in $\theta$,
$\big|\ell(\theta_{T};z)-\ell(\theta^{\prime}_{T};z)\big|\ \leq\ L\,\|\theta_{T}-\theta^{\prime}_{T}\|.$
Taking expectation over all randomness (sample replacement, SGD sampling, and possibly algorithmic self-mod randomness) yields a uniform stability parameter
$\epsilon_{\mathrm{stab}}\ :=\ \sup_{z}\ \mathbb{E}\big[\,|\ell(\theta_{T};z)-\ell(\theta^{\prime}_{T};z)|\,\big]\ \leq\ \frac{C}{m}\sum_{t=1}^{T}\eta_{t}$
with $C:=2LGe^{L}$ (or any problem-dependent constant absorbing smoothness/diameter factors).
Generalization gap.
By standard stability-to-generalization transfer (uniform stability implies $\mathbb{E}[R(\hat{h})-\widehat{R}_{S}(\hat{h})]\leq\epsilon_{\mathrm{stab}}$),
$\mathbb{E}\big[R(\hat{h})-\widehat{R}_{S}(\hat{h})\big]\ \leq\ \frac{C}{m}\sum_{t=1}^{T}\eta_{t}.$
This proves the claim. The bound extends to self-modified schedules because only the realized step sizes $\{\eta_{t}\}$ enter the derivation; any randomness/decision logic is handled by the outer expectation, and the nonexpansive/smoothness arguments hold stepwise. ∎
Meta-policy corollary.
If a metacognitive rule enforces $M_{T}=\sum_{t}\eta_{t}\leq B(m)$, then $\mathbb{E}[\mathrm{gap}]\leq(C/m)\,B(m)$. Choosing $B(m)=\tilde{O}(1)$ gives a $\tilde{O}(1/m)$ expected gap; more generally, $B(m)=\tilde{O}(\sqrt{m})$ gives $\tilde{O}(1/\sqrt{m})$, etc.
15 Full Proofs for Substrate Self-Modification
Assumptions.
i.i.d. data; loss in $[0,1]$; ERM/AERM as specified. A “substrate” is a computational model hosting the same specification $(H,Z,A)$ unless explicitly stated. “CT-equivalent” means mutually simulable (e.g., universal TM, RAM, $\lambda$-calc) with finite simulation overhead independent of $m$.
15.1 CT-invariance (Thm. 8)
Proof.
Let $\mathcal{P}$ be a learning problem that is distribution-freely PAC-learnable on $F$ by algorithm $\mathsf{Alg}$ producing $\hat{h}(S)\in H$ with sample complexity $m(\epsilon,\delta)$. Since $F^{\prime}$ is CT-equivalent to $F$, there exists a simulator $\mathsf{Sim}_{F^{\prime}\leftarrow F}$ that, given the code of $\mathsf{Alg}$, simulates it on $F^{\prime}$ with bounded overhead independent of the sample size $m$. Running $\mathsf{Sim}_{F^{\prime}\leftarrow F}(\mathsf{Alg})$ on $F^{\prime}$ yields the same output hypothesis $\hat{h}(S)$ for any dataset $S$. Hence the distributional correctness (and therefore sample complexity) is unchanged. Computation time may increase by a constant/ polynomial factor but PAC definitions are insensitive to runtime. Therefore $\mathcal{P}$ remains distribution-freely PAC-learnable on $F^{\prime}$ with the same $m(\epsilon,\delta)$ up to constants. ∎∎
15.2 Finite-state downgrade impossibility (Prop. 8)
Modeling the downgrade.
A “finite-state substrate” has a fixed number $N$ of persistent states (independent of $m$) available to the learner across training and prediction. Internally, any learning procedure is a transducer whose internal memory is one of $N$ states; weights/parameters cannot grow with $m$.
Problem and intuition.
Consider thresholds on $\mathbb{N}$ (or $\{0,1\}^{*}$ interpreted as integers): for $k\in\mathbb{N}$, define
$h_{k}(x)=\mathbb{1}\{x\geq k\}.$
[…]
Lemma (Pumping-style indistinguishability).
For any finite-state learner $\mathsf{L}$ with $N$ states and any $m>N$, there exist two training samples $S,S^{\prime}$ of size $m$ such that: (i) $\mathsf{L}$ ends in the same internal state on $S$ and $S^{\prime}$ (with the same output hypothesis), yet (ii) there exists a threshold target $h_{k}$ and a test point $x^{\star}$ on which the two training histories require different predictions to achieve small risk.
Proof.
Process the sorted stream of distinct integers $\{1,2,\dots\}$; after $N$ distinct “milestones” the finite machine must revisit a state (pigeonhole principle). Construct $S$ and $S^{\prime}$ that are identical except for the placement of one positive/negative label beyond the $N$th milestone so that the same terminal state is reached but the correct threshold differs (choose $k$ between the conflicting milestones). Then a test point $x^{\star}$ between those milestones is labelled differently by the two Bayes-consistent thresholds. Since the learner outputs the same hypothesis after $S$ and $S^{\prime}$, it must err on one of the two underlying distributions by a constant amount (bounded away from 0).
Proof of Prop. 8.
[…]
Remark (other families).
An identical argument works for parities over $\{0,1\}^{d}$ with unbounded $d$ or for intervals on $\mathbb{N}$, and more generally any family whose Bayes-consistent parameter cannot be encoded in $O(1)$ bits independent of $m$.
15.3 Beyond-CT substrates (Prop. 8)
Proof.
If $F^{\dagger}$ is stronger than Turing but the measurable hypothesis family under consideration remains $H$, then PAC learnability—defined only in terms of approximation and sample complexity—depends on $H$ and the data distribution, not on computational strength. Thus the sample complexity guarantees are unchanged (computation of ERM/AERM might become easier, but that is outside classical PAC). If, however, the substrate edit enlarges the effective family to $H^{\dagger}$ (e.g., enabling hypercomputable predictors or oracular label access), then learnability is governed by the induced policy-reachable family $\mathcal{H}^{\dagger}_{\mathrm{reach}}(u)$, and all results from §4 apply: preservation iff $\sup_{H^{\prime}\in\mathcal{H}^{\dagger}_{\mathrm{reach}}(u)}\text{VC}(H^{\prime})<\infty$, and destruction otherwise. ∎∎
16 Gödel Machine Foundations
Schmidhuber’s Gödel Machine (Schmidhuber,, 2005) provides a theoretical framework for self-modifying agents that make provably optimal self-improvements. We summarize the key components.
16.1 Architecture and Decision Rule
A Gödel Machine consists of an initial problem-solving policy $e(1)$, a proof searcher that tests proof techniques (programs generating formal proofs), an axiomatic system $\mathcal{A}$ encoding hardware semantics, initial software $p(1)$, environmental assumptions, and a utility function $u$, plus a candidate modification program switchprog.
The machine executes switchprog (which may rewrite any part of the system, including the proof searcher itself) only when it has proven a target theorem:
$u\!\big[s(t_{1})\oplus(\texttt{switchbit}(t_{1})=\text{`}1\text{'}),\text{Env}(t_{1})\big]\ >\ u\!\big[s(t_{1})\oplus(\texttt{switchbit}(t_{1})=\text{`}0\text{'}),\text{Env}(t_{1})\big],$ (4)
where $u$ is the utility function, $s(t)$ is system state at time $t$, and $s(t_{1})\oplus(\texttt{switchbit}(t_{1})=b)$ denotes the hypothetical state with $\texttt{switchbit}:=b$. The left side represents executing the modification; the right side represents continuing to search for alternatives.
16.2 Global Optimality
Theorem (Schmidhuber 2005).
Given any formalizable utility function $u$ and assuming consistency of the axiomatic system $\mathcal{A}$, any self-modification obtained through proof of a target theorem is globally optimal: the utility of executing the present switchprog is higher than the utility of waiting for the proof searcher to produce an alternative switchprog later.
16.3 Bias-Optimal Proof Search
The initial proof searcher uses Bias-Optimal Proof Search (BIOPS), allocating time to proof techniques $w$ proportional to their prior probability $P(w)$.
Theorem (Schmidhuber 2005).
If some unknown proof technique $w$ requires at most $f(k)$ steps to produce a proof of difficulty measure $k$, then BIOPS requires at most $O(f(k))$ steps.
Relation to our framework.
While Gödel Machines establish decision-theoretic optimality through formal proof, our work specializes this framework to provide distribution-free PAC learning guarantees by imposing tractable capacity bounds on the reachable hypothesis family.
17 Experimental Details
Dataset.
We generate binary labels from a smooth ground-truth score on $X\in\mathbb{R}^{d}$ (Gaussian inputs with mild interactions), passed through a logistic link, plus bounded feature noise and a small independent label-flip rate to induce a nonzero Bayes error. A single draw is split once into train/validation/test and reused across sequential edits.
Representational axis $\mathcal{M}_{H}$.
We simulate representational self-modification by increasing polynomial degree $k=0,1,2,\dots$ in a logistic model. Each candidate is trained on the fixed train set and accepted or rejected by one of four edit policies (Two-Gate, dest_val_nocap, dest_val, dest_train). We report the test 011 loss of the last accepted hypothesis at each step; curves end when an edit is rejected.
Algorithmic axis $\mathcal{M}_{A}$.
We hold the hypothesis class fixed and continue stochastic logistic training while tracking the cumulative step-mass $M_{T}=\sum_{t}\eta_{t}$. We compare a capped TwoGate policy (halt when $M_{T}$ reaches a preset budget) to a destructive policy. We plot the generalization gap (test-train loss) versus $M_{T}$.
17.1 Experiment Hyperparameters
Table 2: Experimental parameters for representational self-modification.
Parameter Value
Train/Val/Test 150/60/1k
Max Degree 30
Noise ($\sigma$) 1.2
Label Flip 35%
L2 Reg. ($C$) 1.0
Capacity ($K$) 31
VC Const. ($c_{0}$) 0.10
Val. Margin ($\tau_{\text{mult}}$) 0.20
Confidence ($\delta_{V}$) 0.05
Seeds 5
Table 3: Experimental parameters for algorithmic self-modification.
Parameter Value
Task Setup Degree 5 logistic
Train Samples 500
Val/Test 1k/2k
Seeds 20
Max Iter. ($T$) 50,000
Learn Rate ($\eta_{0}$) 0.01
Batch Size 32
L2 Reg. $10^{-5}$
Noise ($\sigma$) 0.6
Label Flip 20%
Compared Policies for $\mathcal{M}_{H}$
  1. 1.
    Two-Gate: The safe policy enforcing both a capacity gate and a validation margin.
  2. 2.
    Destructive: Naively accepts any capacity increase if training loss does not increase.
Compared Policies for $\mathcal{M}_{A}$
  1. 1.
    TwoGate: The stability meta-policy. Training is halted when the cumulative step-mass ($M_{T}=\sum\eta_{t}$) exceeds a budget of $B=2.5$.
  2. 2.
    Destructive: The baseline policy. Training proceeds for the full 50,000 iterations with no constraint on step-mass.